OpenStudy (anonymous):

AGE PROBLEM! PLEASE HELP! Jam is 2 years younger than Sam. 4 years ago, 1/4 the age of Sam was a third of Jam's age. How old is each boy now? So I have working equations but the answer is fraction so I need help :) Let x be the age of Jam Let y be the age of Sam x+2=y y-12x=-44

OpenStudy (anonymous):

i am not sure

OpenStudy (anonymous):

The 2nd equation should look like this: \[\frac{ y-4 }{ 4 }=\frac{ x-4 }{ 3 }\]

OpenStudy (anonymous):

You need to consider, that it is 4 years ago that that counted. Therefore you need to minus y and x with 4.

OpenStudy (alekos):

I don't get a fractional answer

OpenStudy (alekos):

what answer are you given yoyu?

OpenStudy (alekos):

Just use your 1st equation to solve the 2nd equation given by Tommy and you'll have the answer

OpenStudy (anonymous):

oh okay thanks. i'll try again :)

OpenStudy (anonymous):

so the age of jam is 10 and the age of sam is 12? am i right?

OpenStudy (triciaal):